OpenStudy (anonymous):

if a line has the points of 5,9 and 0,-1 what is the slope-intercept equation?

OpenStudy (texaschic101):

first lets find the slope... slope(m) = (y2 - y1)/ (x2 - x1) set 1 (5,9) set 2 (0,-1) m = (-1 - 9)/(0 - 5) m = -10/-5 = 2 < -- slope is 2 y = mx + b using slope 2 and either of the points ...I am using (5,9) 9 = 2(5) + b 9 = 10 + b 9 - 10 = b -1 = b < -- y intercept So your equation is : y = 2x - 1